OpenStudy (girlgoyleh.):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3?

OpenStudy (anonymous):

what is point slope form first?

OpenStudy (anonymous):

y - y_1 = m(x - x_1). plug and chug theh slope and points into the equatoin: http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (mathmate):

Point-slope form of equation of a line Case : Line through a point with known slope \(\color{teal}{m} \)= slope \((\color{blue}{x_0},\color{purple}{y_0})\) = given point Equation of line \(\boxed{\color{red}{L: y-\color{purple}{y_0} = \color{teal}{m} (x-\color{blue}{x_0})}}\) Example: Find equation of the line with slope 2.5 passing through the point (4,3) \(\color{teal}{m} = \color{teal}{2.5}\) \((\color{blue}{x_0},\color{purple}{y_0})=(\color{blue}{4},\color{purple}{3})\) \(L: (y-\color{purple}{3}) = \color{teal}{2.5} (x-\color{blue}{4}) => y-\color{purple}{3}=\color{teal}{2.5}x-10 => y = 2.5x-7\)

OpenStudy (jmark):

Given slope m=-3 (x0,y0)=(1,4) slope equation y-y0=m(x-x0) y-4=-3(x-1) y-4=-3x+3 y=-3x+7 Follow us at: https://plus.google.com/+EduriteScoreMore/posts